OpenStudy (anonymous):

I am sure there is someone here that can easily tackle this problem! Factor: 3x^(-1/2)+9x^91/2)-81^(-3/2) I panic when I see negative fractional exponents! And my factoring skills are weak at best If you can explain any of the steps or the way you think about the problem I would greatly appreciate the advice.

OpenStudy (anonymous):

\[ 3x^{\frac{-1}{2}}+9x^{\frac{1}{2}}-81^{\frac{-3}{2}}\] like this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I was thinking that an x^(1/2) If 81x^(-3/2) could be rewritten as 81x(-3(1/2)) and 3x(-1/2) could be rewritten as 3x^(-1(1/2))

OpenStudy (anonymous):

i can write out what this means without negative exponents, but i cannot see how to factor it

OpenStudy (anonymous):

oops that an x^(1/2) could be factored out *

OpenStudy (anonymous):

oh wait, is there an x in the last term?

OpenStudy (anonymous):

yes there was

OpenStudy (anonymous):

\[3x^{\frac{-1}{2}}+9x^{\frac{1}{2}}-81x^{\frac{-3}{2}}\]

OpenStudy (anonymous):

oh yikes I overlooked that

OpenStudy (anonymous):

good catch. I looked over it so quickly the first time I missed the omission.

OpenStudy (anonymous):

ok that is a whole different story. first off you can factor a 3 out of each term, then also perhaps an \[x^{\frac{1}{2}}\] if you like. lets try it

OpenStudy (anonymous):

Great I get coefficients of 1 3 and 27

OpenStudy (anonymous):

\[3x^{\frac{1}{2}}(x^{-1}+3+27x^{-2})\]

OpenStudy (anonymous):

now i wonder if we can go further...

OpenStudy (anonymous):

cool that exactly where I am at the moment as well

OpenStudy (anonymous):

oh I had -27^(-3)

OpenStudy (anonymous):

no i think it is \[27x^{-2}\]

OpenStudy (anonymous):

because \[\frac{-3}{2}-\frac{1}{2}=-2\]

OpenStudy (anonymous):

i cannot see how to factor the second part, so i think you are done at that step

OpenStudy (anonymous):

ok so the exponents are not being multiplied in this case

OpenStudy (anonymous):

I was thinking it was -27x^(-3(1/2))

OpenStudy (anonymous):

sorry i was off by a minus sign, should be \[-27x^{-2}\]

OpenStudy (anonymous):

when you multiply you add the exponents

OpenStudy (anonymous):

so \[3x^{\frac{1}{2}}\times -27x^{-2}=-81x^{\frac{1}{2}+(-2)}=-81x^{-\frac{3}{2}}\] which is what you want

OpenStudy (anonymous):

ok I think I see how that works so x^(1/2)x^-2(2/2) = x^(1/2)x^(-4/4) or x^(-3/2)

OpenStudy (anonymous):

Thank you so much for your help! This really helped clarify the concept for me.

OpenStudy (anonymous):

you don't need to change the denominator is 2

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

oops meant (-2(2/2) = (-4/2)

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

awesome!